What is the effective monthly interest rate which is equivalent to a quoted rate of 7.9% with semiannual compounding? A) 7.773%. B) 1.031%. C) 0.648%. D) 0.627%. E) 0.612%.
Answer #1

APR = 7.90% semi-annually

EAR = (1 + APR/n)n - 1

EAR = (1 + 0.079/2)2 - 1 = 8.056%

Calculating Monthly Rate,

Monthly Rate(r) = 12[(1.08056)1/12 - 1]

Monthly Rate(r) = 7.773%